Let f(x) = 1-2x g(x)= 3x^2-2
find f(g(x))

Answer by rapaljer(4671)

f(x)=1-2x, g(x)=3x^2-2

f(g(x)) means to take the inside formula g(x) and substitute it in place of the x in the f(x) formula.

f(x)=1-2x
f(3x^2-2)=1-2(3x^2-2)
=1-6x^2+4
=5-6x^2

Sometimes the work simplifying the result can be difficult, sometimes it is not. If you need to see additional explanation on this and examples that are solved IN COLOR, please see my own website. You can click on my tutor name "rapaljer" anywhere in algebra.com. Then click on the link to my website that is given there.

From my Homepage, click on the link "MATH IN LIVING COLOR." Select "College Algebra" and look in Chapter 2. Look for "Section 2.08 Algebra of Functions."

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus